Multiply using FOIL (x+2) (x-5)
Answer by ankor@dixie-net.com(22740)

Multiply using FOIL (x+2) (x-5)
:
First: x * x = x^2
Outer: x * -5 = -5x
Inner: 2 * x = 2x
Last: 2 * -5 = -10
Put it altogether:
x^2 - 5x + 2x - 10
Combine like terms
x^2 - 3x - 10
:
That's how you "FOIL", pretty easy, right?